OpenStudy (anonymous):

How do you graph y=-1/4x

OpenStudy (anonymous):

the line starts at (0 , 0)

OpenStudy (anonymous):

Woops i didnt mean to post it so much lol, what do you mean by making up points?

OpenStudy (anonymous):

sub x=0 y=-1/4(0)=0 plot (0,0) sub x=1 y=-1/4(1)=-1/4 plot(1,-1/4) draw a line right through the points